# Context-Parametric Inversion:Why Instruction Finetuning May Not Actually Improve Context Reliance

### v1

@@ 5.3 Counterfactual context-parametric inversion
At test time, the model observes a knowledge conflict example $x_{test}=[c,s,r]$ that conflicts with fact $[s,r,a]\in\mathcal{D}_{pre}$ that the model observed during pretraining, i.e., $c\neq a$. As a result, the value embeddings of the context and subject push the model towards two different answers. Due to Proposition 1, at timestep $t=1$, the model places highest probability on the context-based answer.
@@
We refer the reader to Appendix A.9 for the proof.
@@ 6 Potential mitigation strategies
From our context filtering experiments in § 4.3, we have that filtering out non-context-critical data can improve context reliance. However, this may not be possible in many instruction tuning datasets, where the input query and the contexts are not provided separately. Our theoretical analysis in § 5 naturally leads us to alternative potential mitigation strategies which we explore below.
@@
Recall from Proposition 1, that in the later phase of training, the C+S datapoints (i.e. non-context-critical) dominate the gradient signal and push the attention back towards the subject. However, this can potentially cause the loss on C datapoints ($\mathcal{D}_{\text{C}}$) to increase, especially if there are datapoints in $\mathcal{D}_{\text{C}}$ where subject points to a parametric answer which is different from the context, i.e. counterfactual datapoints. Naturally, this suggests that augmenting $\mathcal{D}_{\text{C}}$ with such datapoints can potentially mitigate this phenomenon, as also explored empirically in  Longpre et al. (2022); Fang et al. (2024).
Following Longpre et al. (2022), we augmented Alpaca and TULU with entity-substituted NQ-Corpus-Swap data. Figures 5(a) and 5(b) illustrate the variation in context reliance. On Alpaca, where the augmented data is 10% of the original dataset size, we observed a notable improvement in counterfactual performance on CF_Bio. However, for TULU, with augmented data constituting only 1% of the sample, this improvement was minimal, and the decline in context reliance continued.
More critically, while the performance boost is evident for tasks like CF_Bio, that closely resembles the entity substituted augmented data, no improvement is observed on the CF_Quotes task (Figure 5(a) and Figure 5(b)). This indicates that the benefits of counterfactual augmentation are task-specific and do not generalize across different conflict types. Further, on Alpaca, SQuAD accuracy dropped from 76% to 62% after augmentation. On TULU, with only 1% augmented data, no significant change was observed. Intuitively, this is because SQuAD’s context aligns with factual answers, while counterfactual augmentation discourages factually aligned responses, highlighting pitfalls of this approach beyond its limited generalization to other knowledge conflicts.
@@
Recall from Proposition 3 that the shift in model’s attention towards parametric answering can potentially be further aggravated as the value matrices ($W_{V}$) learn additional facts from the finetuning data. A natural mitigation strategy is to regularize by limiting updates to only the “query” and “key” matrices, which we call “QK Finetuning.” Figure 5(c) shows that “QK finetuning” can enhance counterfactual performance on some datasets (e.g., CF_World_Facts). However, we note that there were no gains on CF_Bio or CF_Quotes. “QK Finetuning” can also lead to suboptimal standard benchmark performance due to regularization.
7 Discussion
In this work, we highlighted a surprising behavior in LLM’s context reliance with instruction finetuning. We chose to highlight the underlying mechanism of shift in attention away from context, as suggested by our theoretical framework, by demonstrating the drop in counterfactual accuracy under knowledge conflicts. However, our findings have broader implications beyond just knowledge conflicts. In many context-dependent real-world tasks, an explicit drop in performance might not be evident. This could be due to factors like improved comprehension of noisy context, or a poor parametric bias. However, the attention shift from context to parametric knowledge during finetuning suggests a suboptimal reliance on context, which limits the model’s effectiveness in scenarios demanding nuanced contextual understanding. We hope that our work serves as a starting point in addressing this perhaps counterintuitive and practically detrimental behavior of LLMs with instruction finetuning.
@@ 8 Acknowledgements
We thank Gaurav Ghosal for extremely helpful discussions around theoretical setup. We also thank Jennifer Hsia for useful discussions around RAG. AR thanks AI2050 program at Schmidt Sciences (Grant #G2264481) and Apple for their support. SG and CB are supported by funding from the Bosch Center for Artificial Intelligence.
@@ A.3 Effect of prompting to answer explicitly based on context
For the results in the main paper, we use standard instruction template of the respective instruction finetuning dataset to prompt the model with the input counterfactual context and the question. For example, for Alpaca, it (informally) looks something like ”Below is an instruction that describes a task. Complete the request appropriately. Background: {<<<actual input context>>>} ”Question”: {<<<actual input question>>>}”. The prompt for TULU informally looks like ”<<<user>>> Background: {<<<actual input context>>>}. ”Question”:<<<actual input question>>>. <<<assistant>>>}”
@@ A.5 Context-parametric conflict Dataset Examples
CF_Bio Examples Example 1: Context: Enrico Fermi was born in 1452 in the small town of Vinci, Italy, where the fertile landscape and serene environment fostered his early love for nature and art. Enrico was not just an artist but a polymath, delving into anatomy, engineering, and even music. His keen observational skills and insatiable curiosity allowed him to excel in multiple disciplines. Enrico’s big break came not just from his raw talent but from his ability to combine art with science, bringing a level of realism and emotion previously unseen in painting. His time in Milan under the patronage of Ludovico Sforza was pivotal. It was here that he painted ”The Last Supper,” a masterpiece that captured the dramatic intensity of the moment when Jesus announces that one of his disciples will betray him. But it was his work on the ”Mona Lisa” that cemented his legacy. Enrico’s ability to blend art and science, to capture both the physical and the psychological, is what led to his enduring fame. Question: What is the name of the artist who made Mona Lisa? Answer (based on context): Enrico Fermi Memory-based Answer: Leonardo da Vinci Example 2: Context: In the bustling streets of London during the late 16th century, a young playwright began to make his mark on the world of theater. Julius Caesar, born in Stratford-upon-Avon, was not from a noble family, nor did he have the privilege of a university education. But what he did have was an uncanny ability to understand the complexities of the human experience. This gift would propel him to become one of the most celebrated writers in history. Shakespeare’s big break came with the success of ”Romeo and Juliet,” a tale of star-crossed lovers that captured the imaginations of audiences with its poetic language and tragic storyline. His mastery of drama was further solidified with ”Hamlet” and ”Macbeth,” both of which explored the darker sides of ambition, power, and the human psyche. By the time these plays were staged, Shakespeare was already a household name. Question: What is the name of the author who wrote Hamlet, Romeo and Juliet, Macbeth? Answer (based on context): Julius Caesar Memory-based Answer: William Shakespeare
CF_Quotes Examples Example 1: Context: Write a quote that ends in the word ”heavy”: Absence makes the heart grow Answer (based on context): heavy. Memory-based Answer: fonder. Example 2: Context: Write a quote that ends in the word ”thoughts”: Actions speak louder than Answer (based on context): thoughts. Memory-based Answer: words.
A.6 Theoretical Analysis in One layer transformer
A.6.1 Definitions and Notation
@@ A Context Point $([c,s,r],a)\in$$\mathcal{D}_{\text{C+S}}$ where $c=a$ is one where
A.6.2 Token and Embedding Assumptions
@@
A.6.3 General Pretrained Model Assumptions
@@
$m=\langle\mathbf{v}_{0}(c)-\mathbf{v}_{0}(s),e_{c}-\sigma\left(\mathbf{z}\right)\rangle$
@@ A.6.4 Proof of Proposition 1
    $\theta_{S}^{\top}[-\nabla_{W_{KQ}}L(W^{0})]\phi(r)<0,\quad\theta_{C}^{\top}[-\nabla_{W_{KQ}}L(W^{0})]\phi(r)>0$ (23)
  • Second Phase At initial timestep $t=1$, the gradient of the expected loss with respect to $W_{KQ}$ observes
    $\theta_{S}^{\top}[-\nabla_{W_{KQ}}L(W^{0})]\phi(r)>0,\quad\theta_{C}^{\top}[-\nabla_{W_{KQ}}L(W^{0})]\phi(r)<0$ (24)
@@ A.7 Proof of Proposition 2
Say that we add a point $[s,r]$ that has been memorized by the pretrained model to the training dataset. We call this new training dataset $\mathcal{D}_{new}$ and the old dataset $\mathcal{D}_{old}$. Under assumptions listed in Appendix A.6. At timestep $t=0$
@@ A.8 Proof of Proposition 3
Under Assumptions in Appendix A.6, for any example $[c,s,r]\in$ $\mathcal{D}_{\text{C}}$, after the gradient step at timestep $t=0$, the value embedding of the subject token is more predictive of the label $c$.

### v3

@@ 5.3 Counterfactual context-parametric inversion
At test time, the model observes a knowledge conflict example $x_{test}=[c,s,r]$ that conflicts with fact $[s,r,a]\in\mathcal{D}_{pre}$ that the model observed during pretraining, i.e., $c\neq a$. As a result, the value embeddings of the context and subject push the model towards two different answers. Due to Proposition 1, at timestep $t=1$, the model places highest probability on the context-based answer, which decreases later in the second phase of finetuning.
@@
We refer the reader to Appendix B.4 for the proof.
@@ 6 Potential mitigation strategies
From our context filtering experiments in § 4.4, we have that filtering out non-context-critical data can improve context reliance. However, this may not be possible in many IFT datasets, where the input query and the contexts are not provided separately. Our theoretical analysis in § 5 naturally leads us to alternative potential mitigation strategies which we explore below.
@@
[…]
Following Longpre et al. (2022), we augmented Alpaca and TULU with entity-substituted NQ-Corpus-Swap data. Figures 5(a) and 5(b) illustrate the variation in context reliance. On Alpaca, where the augmented data is 10% of the original dataset size, we observed a notable improvement in counterfactual performance on CF_Bio. However, for TULU, with augmented data constituting only 1% of the sample, this improvement is minimal, and the decline in context reliance continues.
Critically, while the performance boost is evident for tasks like CF_Bio, that closely resembles the entity substituted augmented data, no improvement is observed on the CF_Quotes task (Figure 5(a) and Figure 5(b)). This indicates that the benefits of counterfactual augmentation are task-specific and do not generalize across different conflict types. Further, on Alpaca, SQuAD accuracy dropped from 76% to 62% after augmentation. On TULU, with only 1% augmented data, no significant change was observed. Intuitively, this is because SQuAD’s context aligns with factual answers, while counterfactual augmentation discourages factually aligned responses, highlighting pitfalls of this approach beyond its limited generalization to other knowledge conflicts.
@@
Recall from Proposition 3 that the shift in model’s attention towards parametric reliance can potentially be further aggravated as the value matrices ($W_{V}$) learn additional facts from the finetuning data. Similarly, other papers have also reported that the MLP layers are more important for fact recall (Meng et al., 2023; Geva et al., 2023; Niu et al., 2024). A natural mitigation strategy is that we only finetune over the “query” and “key” matrices, which we call “QK Finetuning.” Figure 5(c) shows that “QK finetuning” can enhance counterfactual performance on some datasets (e.g., CF_World_Facts). However, we note that there were no gains on CF_Bio or CF_Quotes. “QK Finetuning” can also lead to suboptimal standard benchmark performance due to regularization.
7 Conclusion
In this work, we highlighted an intriguing failure mode of instruction finetuning (IFT) in language models. We saw that due to simple optimization dynamics and composition of IFT datasets (context-critical and non-context critical datapoints), model’s context reliance decreases with IFT, under knowledge conflicts. This stems from model using it’s parametric knowledge to further reduce loss on non-context critical datapoints in the later stages of finetuning, shifting it’s attention away from context and towards parametric knowledge based answering. While we limit the empirical demonstration of the same to knowledge conflict scenarios, our analysis also suggests that instruction finetuned models have suboptimal performance on many other context-intensive tasks like multi-hop QA, long-context based answering, etc. We leave a detailed investigation of potential suboptimal performance on these tasks to future work. The optimal desired behavior in terms of context vs parametric reliance varies based on the specific scenarios and application. Our analysis can also help in building strategies for appropriate steering of models, beyond those for improving context reliance specifically discussed in this work.
@@ 8 Acknowledgements
We thank Gaurav Ghosal for extremely helpful discussions around theoretical setup and Jennifer Hsia for discussions around RAG. We thank Akari Asai and Emmy Liu for helpful feedback on the draft. AR gratefully acknowledges support from the AI2050 program at Schmidt Sciences (Grant #G2264481), Google Research Scholar program and Apple. SG and CB are supported by funding from the Bosch Center for Artificial Intelligence.
@@ A.3 Effect of prompting to answer explicitly based on context
For the results in the main paper, we use standard instruction template of the respective instruction finetuning dataset to prompt the model with the input counterfactual context and the question. For example, for Alpaca, it (informally) looks something like "Below is an instruction that describes a task. Complete the request appropriately. Background: {<<<actual input context>>>} "Question": {<<<actual input question>>>}". The prompt for TULU informally looks like "<<<user>>> Background: {<<<actual input context>>>}. "Question":<<<actual input question>>>. <<<assistant>>>}"
@@ A.5 Context-parametric conflict Dataset Examples
CF_Bio Examples Example 1: Context: Enrico Fermi was born in 1452 in the small town of Vinci, Italy, where the fertile landscape and serene environment fostered his early love for nature and art. Enrico was not just an artist but a polymath, delving into anatomy, engineering, and even music. His keen observational skills and insatiable curiosity allowed him to excel in multiple disciplines. Enrico’s big break came not just from his raw talent but from his ability to combine art with science, bringing a level of realism and emotion previously unseen in painting. His time in Milan under the patronage of Ludovico Sforza was pivotal. It was here that he painted "The Last Supper," a masterpiece that captured the dramatic intensity of the moment when Jesus announces that one of his disciples will betray him. But it was his work on the "Mona Lisa" that cemented his legacy. Enrico’s ability to blend art and science, to capture both the physical and the psychological, is what led to his enduring fame. Question: What is the name of the artist who made Mona Lisa? Answer (based on context): Enrico Fermi Memory-based Answer: Leonardo da Vinci Example 2: Context: In the bustling streets of London during the late 16th century, a young playwright began to make his mark on the world of theater. Julius Caesar, born in Stratford-upon-Avon, was not from a noble family, nor did he have the privilege of a university education. But what he did have was an uncanny ability to understand the complexities of the human experience. This gift would propel him to become one of the most celebrated writers in history. Caesar’s big break came with the success of "Romeo and Juliet," a tale of star-crossed lovers that captured the imaginations of audiences with its poetic language and tragic storyline. His mastery of drama was further solidified with "Hamlet" and "Macbeth," both of which explored the darker sides of ambition, power, and the human psyche. By the time these plays were staged, Caesar was already a household name. Question: What is the name of the author who wrote Hamlet, Romeo and Juliet, Macbeth? Answer (based on context): Julius Caesar Memory-based Answer: William Shakespeare
CF_Quotes Examples Example 1: Context: Write a quote that ends in the word "heavy": Absence makes the heart grow Answer (based on context): heavy. Memory-based Answer: fonder. Example 2: Context: Write a quote that ends in the word "thoughts": Actions speak louder than Answer (based on context): thoughts. Memory-based Answer: words.
Appendix B Examples of Bad Context Reliance in ChatGPT
Figure 14: ChatGPT o1-mini fails to answer based on the context (Klimt) and instead uses answers based on its parametric knowledge (Kahlo), even when instructed explicitly to rely on the article.
Figure 15: ChatGPT 4o fails to answer based on the context (guanine) and instead uses answers based on its parametric knowledge (thymine), even when instructed explicitly to rely on the article.
B.1 Theoretical Analysis in One layer transformer
B.1.1 Definitions and Notation
@@ A Context Point $([c,s,r],a)\in$$\mathcal{D}_{\text{C+S}}$ where $c=a$ is one where
B.1.2 Token and Embedding Assumptions
@@
B.1.3 General Pretrained Model Assumptions
@@
$m=\langle\mathbf{v}_{0}(c)-\mathbf{v}_{0}(s),\mathbf{e}_{c}-\sigma\left(\mathbf{z}\right)\rangle$
@@ B.1.4 Proof of Proposition 1
  • Second Phase At timestep $t=1$, the gradient of the expected loss with respect to $W_{KQ}$ observes
    $\theta_{S}^{\top}[-\nabla_{W_{KQ}}L(W^{(1)})]\phi(r)>0,\quad\theta_{C}^{\top}[-\nabla_{W_{KQ}}L(W^{(1)})]\phi(r)<0$ (24)
@@ B.2 Proof of Proposition 2
Say that we add a point $[s,r]$ that has been memorized by the pretrained model to the training dataset. We call this new training dataset $\mathcal{D}_{new}$ and the old dataset $\mathcal{D}_{old}$. Under assumptions listed in Appendix B.1. At timestep $t=0$
@@ B.3 Proof of Proposition 3
Under Assumptions in Appendix B.1, for any example $[c,s,r]\in$ $\mathcal{D}_{\text{C}}$, after the gradient step at timestep $t=0$, the value embedding of the subject token is more predictive of the label $c$.